# LLMs for Time Series: an Application for Single Stocks and Statistical Arbitrage

### v1

@@ 2.3.2 Strategy based on the forecast of the fine tuned version of Chronos
Secondly, we used a very naive solution for fine tuning from the pretrained weights which could be a nightmare in practice (Goodfellow et al. (2013)). We trained Chronos which was initiated at the beginning of the backtest with the pretrained weights. The training was realized on a daily basis during the backtest using the available financial market data starting on every day with the weights of the previous day. On every day $d$, we provided as input to Chronos the updated time series available at day $d$ using the the previous 100 days. We test different parameters for $\tau$ the maximal number of steps for the daily training. We also used different values of the parameter $\alpha$ in Eq 1 so that Chronos receives the EMA. Thanks to that feed, Chronos can adapt its weights according to the properties of the financial time series.
The continuous training led to Chronos model to update its weights for each day of the backtest. We then determined the portfolio weights based on the predictions using the fine-tuned weights instead of the pretrained ones. Finnaly we use exactly the same evaluation’s receipe than the one described in section 2.3.1 with the only difference that the pretrained weights were replaced every day by the fine tuned weights determined at that day.
@@
We do not claim that our methodology for fine tuning is optimal, as there are likely better methods to empirically determine an improved approach by controling overfitting and the loss of the pretrained weights through the analyze of the statistics of the eigenvalues derived from the millions of weights (Martin (2019, 2024)). However, this falls outside the scope of our current study. For instance, the drawback of our methodology is that the pretraining weights are gradually forgotten over time, which is not an ideal solution.
2.3.3 Evaluation of other strategies for comparaison
@@
Third, we compare the results of Chronos to those obtained by replicating the CNN Transformers model of Guijarro-Ordonnez et al. (2022) whose number of parameters is only 169 which appears quite small compared with the 11 million of the Chronos one. We also include the results achieved using autoARIMA from the statsforecast package ( https://pypi.org/project/statsforecast/) as it is a standard benchmark in Machine Learning, as well as the short term reversal STR described in Jegadeesh (1990); Jegadeesh and Titman (1993) which is a well-documented market anomaly that was first noted by Fama (1965).
The short term reversal (STR) strategy was derived in Eq 10 with both $\beta=1-\frac{1}{5}$ and $\beta=1-\frac{1}{20}$ using a simple exponential moving average on residual returns $r_{d+1,i}$ at day $d+1$ and single stock $i$ when extracting common factors from the IPCA, PCA or FF. The portfolio of the STR is then derived in Eq 11 with ${\omega^{\zeta}}_{d}$ using the same methodology as above where N is the number of single stocks.
The AutoARIMA was also fit in a continuous way every day during the backtest using the previous 100 days $\times$ N observations and yielded to forecasts ${\mathbf{A}}_{d}$ and the portfolio weights ${\omega}^{\mathbf{A}}_{d}$ were also derived with the same methodology Eq 12.
@@
3 Our empirical results
We observe that the pre-trained Chronos model with $\alpha=0.3$ effectively identifies opportunities in the financial market, achieving a Sharpe ratio above $3.17$ for PCA over a 15-year period, which corresponds to a t-statistic of $12.273.17\sqrt{15}=$ 12.27. However, trading costs are prohibitive, as including a 3 basis point slippage cost per trade results in negative net Sharpe ratios (see Table 1).
@@ 3 Our empirical results
Table 1: Simulation of the Gross sharpe ratio of the strategy based on the zero-shot pretrained prediction of Chronos when using as input the exponential moving average of daily residual returns through using either the IPCA, the PCA or FF. $\alpha$ is the parameter of the EMA from Eq 1. The period is 2002-2016.
FF PCA IPCA
@@
Table 2: Simulation of the Gross sharpe ratio of the strategy $\alpha$ is the parameter of the EMA in Eq 1. $\beta$ is the parameter of the EMA in Eq 10. $\tau$ is ’max steps’ input in the fine-tuned version of Chronos. The period is 2002-2016.
(a) zero-shot pretrained prediction of Chronos. $\alpha=0.3$. The orange is for pca, the green for FF, the blue for ipca.
(b) zero-shot pretrained prediction of Chronos. $\alpha=0$. The orange is for pca, the blue for ipca, the green for FF.
(c) STR $\beta=0.8$. The blue is for ipca, the orange for pca, the green for FF.
@@
Additionally, we note a decline in profitability over time, suggesting that markets may be becoming more efficient or that opportunities are increasingly challenging to capture or that returns used to be more negatively autocorrelated before 2008 (see Figure 1). However, at least until 2007, it was easier for AI to capture inefficiencies.
It is particularly interesting to observe that the pre-trained version with $\alpha=0$ is ineffective until 2007 but seems to work after 2008 (see Figure 1). In our interpretation, Chronos is pre-trained on data where ’trend’ serves as an efficient indicator, whereas in our dataset, residual returns tend to be negatively autocorrelated in the short term. We believe that $\alpha=0.3$ is optimal, as it offsets this effect, helping Chronos to overcome biases from its trend-oriented training dataset.
Setting $\alpha=0.3$ artificially aids Chronos; however, it ultimately gets very correlated to the Short-Term Reverseal (STR) strategy with $\beta=0.3$ but fails to outperform it . In other words, when Chronos is provided with the Exponential Moving Average (EMA), its performance does not exceed that of a zero forecast. Nevertheless, it avoids being affected by detrimental noise, which is already a positive outcome.
Table 2 presents the Sharpe ratios for the fine-tuning case as well as for the benchmarks.
When setting $\alpha=0$, Chronos requires fine-tuning with $\tau=15$ to achieve a Sharpe ratio of $0.24$, which corresponds to a t-statistic of $0.920.24\sqrt{15}=$ 0.92. It appears to work well until 2008 (see Figure 1), but after that, the pre-trained configuration may have been completely forgotten due to the numerous fine-tuning processes performed since 2002. It might be interesting to test a version where there is a regular reinforcement of the pre-trained configuration to ensure it remains in memory.
Additionally, the correlation with the STR strategy is not significant when $\alpha=0$ and $\tau=15$. In contrast, the CNN-Transformer appears to be primarily a linear combination of STR strategies at different time scales. This suggests that the opportunities captured by Chronos may be more complex than those driven by basic mean reversion. We also observe that the autoARIMA model, which is a classical benchmark in Machine Learning underperforms the STR model, demonstrating that fitting a model is particularly challenging when the data are nearly random and contain significant noise. This results in underperformance compared to more rigid models like STR.
We can also see that resizing the weights inversely proportional to the volatility improves the Sharpe ratio for Chronos, as well as for the benchmarks.
Finally, it is interesting to note that the optimal $\tau$ for training appears to be 15. When $\tau=40$, the Sharpe ratio decreases, suggesting that Chronos may lose some of its pre-trained intelligence.
@@ 4 Conclusion
Nevertheless, we believe that specialized models, such as those by Valeyre (2024), which theoretically capture well-established opportunities in an optimal way (like trends), will always prove more efficient, while AI could serve as a valuable tool for identifying more complex opportunities.
That belief is justified by the case of the strong outperformance of the STR compared to AutoARIMA, which can capture more complexity but whose noisy fit makes it suboptimal and overly erratic.
@@ B.2 Parameters of Fine tuned version of Chronos
From the inital version set from the pretrained case at the begining of the period of the backtest, we update the weights of the chronos model at every day of the backtest from the weights obtained at the previous day by executing every day 10 times on 10 subgroups of the universe $\tau$ ”steps” inside ”trainer.train()” ($\tau$ successive corrections of the weights using the ”adamw torch fused” gradient algo) per day in the backtest using the past 100 days. $\tau$ was tested to 5, 15 and 40. $\tau$ is the maximum training steps, i.e. the ”max steps” parameter in the ”TrainingArguments” method. We used the following parameters in python:

### v2

@@ 2.3.2 Fine-tuned version
Secondly, we employed a very naive approach to fine-tuning from the pretrained weights, which can be highly challenging and unstable in practice Goodfellow et al. (2013). Thanks to that feed, Chronos can adapt its weights according to the properties of the financial time series. We trained Chronos which was initiated at the beginning of the backtest with the pretrained weights. The training was realized on a daily basis during the backtest using the available residual returns data starting on every day with the weights of the previous day. On every day $d$, we provided as input to Chronos the updated time series available at day $d$ using the the previous 100 days. We test $5$ and $15$ as values for $\tau$, the maximal number of steps in the iteration process. $\tau$ is a critical parameter of the algorithm, detailed in Appendix B.2, governing the daily training process. If $\tau$ is set too high, Chronos may completely forget its pretrained knowledge, leading to severe overfitting—especially given the relatively small size of our dataset compared to the model’s 11 million parameters. Moreover, even a moderate value such as $\tau=15$ is already highly computationally demanding: it required approximately one week of computation on our setup, which includes 22 Intel(R) Core(TM) i7-14700KF 3.40GHz processors, two NVIDIA GeForce RTX 4060 Ti GPUs with 80 GB of combined VRAM, and 128 GB of RAM. We also used different values of the parameter $\alpha$ in Eq 1 which adjusts the input of Chronos. The continuous training led to Chronos model to update its weights for each day of the backtest leading to a continuous outsample result. We then determined the portfolio weights based on the predictions using the fine-tuned weights instead of the pretrained ones.
@@
We do not claim that our methodology for fine tuning is optimal to empirically determine other approaches by controling overfitting and the loss of the pretrained weights through the analysis of the statistics of the eigenvalues derived from the millions of weights (Martin (2019, 2024)). However, this falls outside the scope of our current study. For instance, the drawback of our methodology is that the pretraining weights are gradually forgotten over time, which is not an ideal solution. We give more detail of the methodology we used for fine tuning in Section B.2 of the Appendix.
We did perform an analysis using WeightWatcher (Martin (2024)) on both the pretrained and fine-tuned models, and found no signs of overfitting or underfitting. The results of this analysis (Fig 2 in Sec B.3 of the Appendix ) which is distribution of the power law coefficient of the eigenvalues distributions layer per layer are included in the appendix and show that the power-law exponents fall within the range of 2 to 6 in the fine-tuned case, indicating no risk of overfitting (which would be suggested by exponents below 2) or underfitting (suggested by exponents above 6). The statistics from the fine-tuned model are even more favorable than those of the pretrained model, which shows some exponents above 6, suggesting a minor risk of underfitting.
2.3.3 Comparison
@@
Third, we compare the results of Chronos to those obtained by replicating the machine learning model with a small number of 169 parameters of Guijarro-Ordonnez et al. (2021), who used a CNN Transformers model. Their number of paramaters is small compared with the 11 million of the Chronos one. They used a CNN Transformers model. We also include the results achieved using standard autoARIMA from the statsforecast package ( https://pypi.org/project/statsforecast/) as it is a standard benchmark in Machine Learning, as well as the Short Term Reversal (STR) described in Jegadeesh (1990); Jegadeesh and Titman (1993) which is a well-documented market anomaly that was first noted by Fama (1965) as the same anomaly is captured by Chronos.
Both AutoARIMA and Short Term Reversal (STR) appear to be well-suited benchmarks, as they capture anomalies that are very similar to those detected by Chronos within our framework. Guijarro-Ordonnez et al. (2021) captures also Short Term Reversal (STR). The returns between their portfolios are correlated.
The Autoregressive integrated moving average (ARIMA) specification is part of the ARMA statistical models that typically predict future values based on explicit past values. The class of ARIMA models has been widely applied through Box and Jenkins (1970) methodology and has proved fruitful for time series modeling according to Engle (1978). They are made with three components: First, an Auto-Regressive (AR) component based on a number of lagged values with order (p). Second, a Moving Average (MA) component based on the number of lagged residual terms. Third, an Integrated (q) component based a number of differencing orders used to make the data stationary on a degree (d). The auto-ARIMA specification use information criteria such like AIC or BIC to determine optimally the ARIMA parameters (p,d,q), where the best ARIMA model for the time series is one with the smallest AIC or BIC. The AutoARIMA was also fit in a continuous way every day during the backtest using the previous 100 days $\times$ N observations and yielded to forecasts ${\mathbf{A}}_{d}$ and the portfolio weights ${\omega}^{\mathbf{A}}_{d}$ were also derived with the same methodology Eq 12.
Short Term Reversal (STR) financial strategy has been a market anomaly that was first observed by Fama (1965) and first documented by Jegadeesh (1990). It belongs to the category of price trend-following strategies in the industry of portfolio management and it captures the Short Term mean-reversion behavior. In practice this dynamic strategy requires sorting stocks into deciles according to their past-month returns in order to buy losers (stocks in the bottom decile) and sell winners (stocks in the top decile). It is a dynamic strategy because it is rebalanced on a monthly basis. It is also a robust and statistically significant strategy because Jegadeesh (1990) unveil a monthly extra return of 2% (see for a detailed discussion, see e.g. Da et al. (2014)). The Short Term Reversal (STR) strategy was implemented using an exponential moving average of residual returns, as defined in Equation 10, with two parameter settings: $\beta=1-\frac{1}{5}$ and $\beta=1-\frac{1}{20}$. The portfolio of the STR is then derived in Eq 11 with ${\omega^{\zeta}}_{d}$, the weights of the portfolios, using the same methodology as for the Chronos forecast where N is the number of single stocks.
@@
2.4 Performance measures
We based our results on the gross Sharpe ratio metric of the portfolio which corresponds to the outsample annualized average of daily returns of the portfolio divided by the annualized standard deviation of daily returns, meaning that we do not consider slippage and trading costs. This metric is standard in the portfolio management literature and can be associated with the t-statistic by multiplying it by the square root of 15, corresponding to the 15 years in our sample period. If the Sharpe ratio of the portfolio is positive and the associated t-statistic exceeds 3, we can consider that the prediction methodology has succeeded in capturing some market anomalies through its forecasts. Fama and French (2015) typically assess anomalies by comparing the performance of portfolios formed on different quantiles (e.g., quintiles) of a given factor. They consider an anomaly to exist if the return spread between extreme portfolios is statistically different from zero, based on a t-statistic. In this sense, our use of the Sharpe ratio—or its associated t-statistic—is closely related to their methodology. The Sharpe ratio is particularly relevant as it provides an aggregate performance measure across the entire universe and is a simple yet widely used metric in asset management to evaluate whether a strategy is efficiently generating profits. A net Sharpe ratio of 1 is already considered excellent from the perspective of a final investor.
3 Empirical results
We can see that resizing the weights of the profolio inversely proportional to the volatility improves the Sharpe ratio for Chronos slighly, as well as for the benchmarks so that we chose to report only the results obtained from this version for all methods.
We observe that the pre-trained Chronos model with the autoregressive coefficient $\alpha=0.3$ effectively identifies opportunities in the financial market (see Table 2), achieving a Sharpe ratio above $3.17$ for PCA over a 15-year period, which corresponds to a t-statistic of $3.17\sqrt{15}=12.27$. However, trading costs are prohibitive, as including a 3 basis point slippage [^3] A transaction cost of 3 basis points per trade is consistent with the average market impact—typically between 2 and 3 basis points for very small orders—and brokerage fees of approximately 1 basis point. With $\alpha=0.3$, the gross Sharpe ratio for PCA is 3.17, but it drops to -1.49 after accounting for a 3 basis point trading cost. cost per trade results in negative net Sharpe ratios. Additionally, we note a decline in profitability over time, suggesting that markets may be becoming more efficient or that opportunities are increasingly challenging to capture or that returns used to be more negatively autocorrelated before 2008 (see Figure 1). However, at least until 2007, it was easier for AI to capture inefficiencies.
@@ 3 Empirical results
Table 2: Simulation of the Gross sharpe ratio of the strategy with the resized version (ie. Eq 5) to reduce risk on volatiles stocks based on the zero-shot pretrained prediction of Chronos when using as input the exponential moving average of daily residual returns through using either the IPCA, the PCA or $F$. $\alpha$ is the parameter of the EMA from Eq 1. The period is 2002-2016.
$F$ PCA IPCA
@@
Table 3: Simulation of the Gross sharpe ratio of different strategies using the resized version (ie. Eq 5) to reduce risk on volatiles stocks. $\alpha$ is the parameter of the EMA in Eq 1. $\beta$ is the parameter of the EMA in Eq 10. $\tau$ is ’max steps’ input in the fine-tuned version of Chronos. The period is 2002-2016.
(a) zero-shot pretrained prediction of Chronos. $\alpha=0.3$. The orange is for pca, the green for $F$, the blue for ipca.
(b) zero-shot pretrained prediction of Chronos. $\alpha=0$. The orange is for pca, the blue for ipca, the green for $F$.
(c) STR $\beta=0.8$. The blue is for ipca, the orange for pca, the green for $F$.
@@
It is particularly interesting to observe that the pre-trained version with $\alpha=0$ is ineffective until 2007 but seems to work after 2008 (see Figure 1). In our interpretation, Chronos is pre-trained on data where ’trend’ serves as an efficient indicator, whereas in our dataset, residual returns tend to be negatively autocorrelated in the short term. We believe that $\alpha=0.3$ is optimal, as it offsets this effect, helping Chronos to overcome biases from its trend-oriented training dataset. Setting $\alpha=0.3$ artificially helps Chronos. However, it ultimately gets very correlated to the Short-Term Reverseal (STR) strategy with $\beta=0.3$ but fails to outperform it . In other words, when Chronos is fed with the Exponential Moving Average (EMA), its performance does not exceed that of a zero forecast. Nevertheless, it avoids being affected by detrimental noise, which is already a positive outcome.
Table 3 presents the Sharpe ratios for the fine-tuning case as well as for the benchmarks. When setting $\alpha=0$, Chronos requires fine-tuning with $\tau=15$, which is the number of iterations of the algorithm, to achieve a Sharpe ratio of $0.24$, which corresponds to a t-statistic of $0.24\sqrt{15}=0.92$. It appears to work well until 2008 (see Figure 1), but after that, the pre-trained configuration may have been completely forgotten due to the numerous fine-tuning processes performed since 2002. It might be interesting to test a version where there is a regular reinforcement of the pre-trained configuration to ensure it remains in memory. Additionally, the correlation with the STR strategy is not significant when $\alpha=0$ and $\tau=15$. In contrast, the CNN-Transformer appears to be primarily a linear combination of STR strategies at different time scales based on its empirical correlation with such a basktet. This suggests that the opportunities captured by Chronos may be more complex than those driven by basic mean reversion process. We also observe that the autoARIMA model, which is a classical benchmark in Machine Learning underperforms the STR model, demonstrating that fitting a model is particularly challenging when the data are nearly random and contain significant noise. This results in underperformance compared to more rigid models like STR. Finally, it is interesting to note that the optimal $\tau$ for training appears to be 15. When $\tau=40$, the Sharpe ratio decreases, suggesting that Chronos may lose some of its pre-trained intelligence.
@@ 4 Conclusion
Nevertheless, the specialized models, which theoretically capture well-established opportunities in an optimal way, will always prove more efficient, while AI could serve as a valuable tool for identifying more complex opportunities. That assertion is supported by the case of the strong outperformance of the STR compared to AutoARIMA, which can capture more complexity but whose noisy fit makes it suboptimal and overly erratic.
@@ B.2 Parameters of Fine tuned version of Chronos
Starting from the initial pretrained weights at the beginning of the backtest period, we updated the Chronos model’s weights daily throughout the backtest. Each day, this update was performed by running the training procedure on 10 subgroups of the asset universe, using the previous 100 days of data. For each subgroup, training was conducted with $\tau$ iterations per day, corresponding to the ”max steps” parameter in the ”TrainingArguments” configuration of the ”trainer.train()” method. In other words, $\tau$ successive weight updates were applied using the ’AdamW Torch Fused’ gradient optimization algorithm. We tested $\tau$ values of 5, 15, and 40. The specific Python training parameters used are listed below:
@@
B.3 Analysis by WeightWatcher
We assessed the quality of both the pretraining and fine-tuning of Chronos using the analysis method described at https://weightwatcher.ai/. Fig 2 shows that the power-law exponents fall within the range of 2 to 6 in the fine-tuned case, indicating no risk of overfitting (which would be suggested by exponents below 2) or underfitting (suggested by exponents above 6). The statistics from the fine-tuned model are even more favorable than those of the pretrained model, which shows some exponents above 6, suggesting a minor risk of underfitting.
Figure 2: Analysis of the Weightwatcher (Martin (2024)) of the distribution of eigenvalues shows that there is no overfitting nor underfitting in both pretrained and finetuned case.